Question Help

A gardener has 37 feet of fencing to be used to enclose a rectangular garden that has a border 2 feet wide surrounding it.
Use this information to answer the following.


​(a) If the length of the garden is to be twice its​ width, what will be the dimensions of the​ garden? 

Answer by Boreal(15235)

length= l+4, 2 feet on either side
width=w+4, same
but l=2w
so dimensions are 2w+4+w+4
the perimeter is twice their sum or 2(3w+8)=6w+16
6w+16=37
6w=21
w=3.5 feet
length is 7 feet
width plus fencing is 7.5 feet and length is 11 feet, because both sides have 2 extra feet.
3.5 x 7 feet answer
perimeter is 7.5 x 11, with perimeter 37'.